# Demystifying Language Model Forgetting with Low-rank Example Associations

### v5

@@ 6 Conclusions
In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning. We showed that simple low rank patterns are dominant in the example associations and compared the complexity of the associations across model types and sizes. We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks. We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweigh examples for replay with predicted forgetting. Future works can extend the study to a continual learning setup where new domains or tasks are sequentially.
Impact Statement
@@
By analyzing and mitigating forgetting, our research tries to demystify and reduce forgetting in LLM fine-tuning. The better understanding alongside the reduced forgetting in turn encourage model developers to promptly update their LLMs to address limitations of the models and improve their models in continuing efforts. We expect the broader application of the continual learning practice will reduce training costs compared to re-training models, and ultimately result in more powerful models under a controlled training cost.
Appendix A Dataset, Model, and LM Training Details
@@
Subsample of upstream dataset. For OLMo experiments, we sample 141,876 text chunks with length 2,048 from Dolma v1.6-sample as upstream examples. For OLMo-7B-Instruct, we randomly sample an approximately balanced number of examples from each task in Tulu, and filter out examples with input length that exceeds 2,048 (the limit of OLMo models) after tokenization. This results in 10,718 examples. For OLMo2, we sample 70,000 text chunks with length 2,048 from OLMo2-Mix. For MPT and Pythia, we sample 10,000 2,048-token text chunks from RedPajama and the Pile respectively.
@@
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 7 and 8 in our experiments.
@@
Training and evaluation details. For full-parameter fine-tuning of non-instruction-tuned LLMs of all types, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$. For LoRA fine-tuning, we set the rank of adapters as 64 in all our experiments and use a rate of $10^{-4}$. We train the models for 625 steps with an effective batch size of 8. For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA, considering the small size of the training sets, we train the models only for 37 steps with an effective batch size of 8. We use HuggingFace Transformers library for training and VLLM library for efficient inference. The statistics of forgetting are obtained in a single run.
@@
Figure 9: A side-by-side comparison between the matrices of forgetting, inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$), and the negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$) we examines in Sec. 3.3.
(a) OLMo-7B-Instruct; binary forgetting over FLAN after full-parameter fine-tuning on unseen instruction-tuning tasks.
(b) OLMo-7B-Instruct (LoRA); binary forgetting over FLAN after full-parameter fine-tuning on unseen instruction-tuning tasks.
@@
Figure 10: Visualized matrices of associations between learned tasks and forgotten examples measured with binary exact match drops.
Appendix B Details of Forgetting Prediction and Replay
@@
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 7 and 8. The train-test split for the in-domain tasks is randomly generated.
@@
Training and evaluation details. We use Surprise Library 1.1.3111https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, MF, and KNN prediction models. For MF, we set the dimension of the learnable features as 5. We train the regression models for 1,000 epochs over the association matrices.
@@
For in-domain test splits, we randomly sample 30 upstream examples and assume the ground truth forgetting is known for these examples. This is required for predicting forgetting on the rest of upstream examples by additive linear, MF, and KNN methods. We repeat the experiment 10 times.
Table 4: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
KNN-Offline 67.49 42.24 74.33 77.07 44.30 80.09 84.91 67.54
@@
We used OLMo-1B models as the trainable example encoders in the implementation of the prediction method by (Jin & Ren, 2024) that relies on inner products of trained example representations. At inference, given an upstream example, we compute the averaged dot-product with all examples in the learned task. We note that at inference time the approach does not require ground truth forgetting of a small number of examples. For a fair comparison with other matrix completion methods, we replace the prediction of the approach with ground truth forgetting on these examples.
(a) OLMo-7B (Full)
(b) OLMo-7B (LoRA)
@@
Figure 11: Reconstruction of $Z$ in OLMo-7B experiments with $k$-th singular value and vectors. Higher values of $k$ capture finer-grained details in $Z$.
@@
Replaying upstream examples in fine-tuning. We sparsely replay 1 mini-batch of 8 upstream examples every 32 steps of model update while fine-tuning on new tasks. Given predicted or ground truth forgetting $z_{i,1..J}$ on upstream examples $x_{1..J}$ when learning a new task $T_{i}$, we sample upstream examples to replay from a categorical distribution where $p(x_{j})\propto\exp(z_{i,j}/\tau)$, where $\tau$ is a temperature hyperparameter set as 0.1. The hyperparameter $\tau$ is tuned on a single validation task by using ground truth forgetting $Z$.
Appendix C Details of the Example Similarity Metrics
@@
Textual similarity. We measure textual cosine similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized features over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$. We also measure text representation similarity with final layer representations of OLMo-1B.
Inner products between projected gradients and model weight updates. The increase of the log perplexity ${z_{ij}}$ can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following (Park et al., 2023; Xia et al., 2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
@@
Appendix D Downstream Task Evaluation with Task-Specific Metrics
We evaluate downstream task performance of LMs before and after fine-tuning on 8 tasks from Dolly with LM-Evaluation-Harness framework (Gao et al., 2024). For OLMo-7B models, we evaluate on the same set of downstream tasks in OLMo technical report (Groeneveld et al., 2024). For OLMo-7B-Instruct models, we evaluate on Open LLM Leaderboard tasks [^2] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about. For fine-tuned models, we compare no replay, replaying random examples, and replaying with forgetting predicted by offline KNN. Tables 4 and 5 summarize the results.
We notice that fine-tuning OLMo-7B on Dolly improves downstream performance on most of the downstream tasks. This aligns well with the purpose of fine-tuning a LM that is not instruction-tuned. Nevertheless, we notice performance degradation on two of the tasks, namely ARC-Easy and Sciq, which indicates forgetting. Although offline KNN achieves higher accuracy scores on these two tasks compared to no-replay (67.48 to 67.34, 84.91 to 84.09), we do not find the improvement statistically significant. For OLMo-7B-Instruct, fine-tuning on Dolly only improves performance on MUSR. The models clearly suffer from forgetting on the other tasks such as IFEval. Offline KNN achieves higher scores than random or no replay on IFEval (18.28 compared to 18.19 and 17.72), but we could not conclude about the significance of the improvement.
To summarize, we do not see clearly reduced forgetting in downstream task performance (evaluated with task-specific metrics) by replaying random or chosen examples. We conjecture that replay-based approaches are not sufficient to mitigate forgetting on their own, and can be combined with other approaches such as careful learning rate scheduling or parameter regularization. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
Table 5: Downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on dolly tasks.
@@ Appendix D Downstream Task Evaluation with Task-Specific Metrics
No Replay 17.51 37.63 17.72 40.95 26.39
Random 17.45 37.63 18.19 40.74 26.46
KNN-Offline 17.45 37.49 18.28 40.74 26.48
@@
Learned Tasks More(+) / Less(-) Forgotten Domain Learned Tasks More(+) / Less (-) Forgotten Domain
@@
$k=2$ dolly/general_qa StackOverflow (-) flan/mnli_mismatched None
@@
Appendix E Towards Interpreting Fine-Grained Associations
We visualize progressive reconstruction with $k$-th singular value and singular vectors for OLMo experiments in Figure 11. The visualization exemplifies complicated associations that is not captured by the simple multiplicative model ($k=1$). For example, on OLMo-7B (LoRA) and when $k=2$, we see a single row and column with significantly larger forgetting than the others.

### v6

@@ 6 Conclusions
In this paper, we empirically analyzed the associations between learned and forgotten examples in LM fine-tuning. We showed that simple low rank patterns are dominant in the example associations and compared the complexity of the associations across model types and sizes. We showed the example associations alone offer useful information to predict example forgetting when fine-tuning LMs on new tasks. We demonstrated the practical utility of our analysis by showing reduced forgetting as we reweigh examples for replay with predicted forgetting. Future works can extend the study to a continual learning setup where new domains or tasks are learned sequentially. We discuss limitations of the work in Appendix A.
Appendix A Limitations
Although our work tries to be extensive, we consider there are limitations that can be taken as future works.
Forgetting under sequential task learning. In this paper, we analyze and predict forgetting of models that are separately fine-tuned over diverse downstream tasks. We did not explore forgetting under continual or sequential task learning. Although we consider time series forecasting to be relevant to this generalized forgetting prediction problem, we decide to leave it as future work.
Theoretical analysis of low-rank example associations. While we conclude that low-rank example associations are prevalent in LLM forgetting from empirical results, we do not have a theoretical guarantee on why example associations must be low rank, and in which cases the example associations become more complicated. We leave a theoretical understanding of low-rank example associations as future work.
Scope of the experiment setups. Our analysis is restricted to supervised fine-tuning of LMs; other learning setups such as reinforcement learning is not studied. Besides, although we try to cover a broad category of newly learned tasks, the coverage may not be extensive. Due to limitations in computational resources, it was also not possible to repeat experiments under extensive hyperparameter configurations (e.g., learning rate schedulers), which we leave for future work. Nevertheless, we provide an analysis of more learning rate and batch size setups in Appendix H.
Requirement of open access to upstream data. Our analysis requires open access to upstream data to evaluate their forgetting. As a result, our analysis is limited to open-data LLMs. In addition, similar to almost every replay-based algorithm, our forgetting mitigation algorithm requires access to upstream data. We hope that our work can inspire researchers that build closed-data LLMs with in-house data.
Scale of the upstream examples. Due to computational resource restrictions, we significantly subsampled upstream pretraining corpora of LLMs in our analysis. Although expanding the set of upstream examples would not theoretically affect the low-rank observations, we decide to leave larger scale analysis as future works. Besides, we consider that matrix completion based forgetting prediction permits sparse association matrices, where upstream example forgetting $z_{ij}$ are sparsely collected, which we will explore in future work.
Separating out intended forgetting from unintended forgetting. Aligning with most prior works on continual learning, our experiments focus on mitigation of forgetting. We consider the goal aligns with our experiment setup, as upstream examples from Dolma and Tulu are carefully curated and cleaned, and the newly learned tasks intend to accumulate knowledge instead of overwrite previous knowledge. However, we highlight that forgetting is not always detrimental; in fact, the models are often expected to forget (unlearn) sensitive, noisy, and outdated examples (Nguyen et al., 2022). We believe that our analysis of forgetting helpful for future works that performs targeted forgetting of knowledge.
Appendix B Broader Impact
@@
Our research tries to better demystify and mitigate forgetting in LLM fine-tuning. We expect that the better understanding alongside the reduced forgetting can, in turn, encourage model developers to promptly update their LLMs to address limitations of the models and improve their models in continuing efforts. We expect the broader application of the continual learning practice will reduce training costs compared to re-training models, and ultimately result in more powerful models under a controlled training budget.
Although we do not see direct negative impact of predicting example forgetting, we highlight that in real-world continual learning setups, blindly mitigating forgetting may result in outdated knowledge and data privacy breaches in LLMs. Dissecting beneficial and intended forgetting from unintended or catastrophic forgetting requires attention in real-world setups.
(a) OLMo-1B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(b) OLMo-7B (LoRA); forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(c) MPT-7B; forgetting over Dolma after full-parameter fine-tuning on FLAN and Tulu.
(d) OLMo-7B-Instruct; forgetting over Tulu after full-parameter fine-tuning on unseen instruction-tuning tasks.
Figure 6: Additional visualized matrices of associations between learned tasks and forgotten examples. We plot forgetting (log-perplexity increase) that occurs on an upstream example (in $x$-axis) after learning a new task (in $y$-axis). Log-perplexity increase can be zero or negative, which indicates no forgetting.
Appendix C Dataset, Model, and LM Training Details
@@
Subsample of upstream datasets. For OLMo experiments, we sample 141,876 text chunks with length 2,048 from Dolma v1.6-sample as upstream examples. For OLMo-7B-Instruct, we randomly sample an approximately balanced number of examples from each task in Tulu, and filter out examples with input length that exceeds 2,048 (the limit of OLMo models) after tokenization. This results in 10,718 examples. For OLMo2, we sample 70,000 text chunks with length 2,048 from OLMo2-Mix. For MPT and Pythia, we sample 10,000 2,048-token text chunks from RedPajama and the Pile respectively.
@@
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 12 and 13. We also annotate tasks used as in-domain training or test tasks.
@@
Training and evaluation details. For full-parameter fine-tuning of non-instruction-tuned LLMs of all types, we train the model for 1,000 steps with an effective batch size of 8 and a linearly decaying learning rate of $2e^{-6}$. The learning rate is chosen among {$1e^{-6}$, $2e^{-6}$, $5e^{-6}$} that achieves best average validation perplexity after fine-tuning OLMo-7B on 5 randomly chosen tasks from FLAN. For OLMo-7B-Instruct and MMLU, BBH, TruthfulQA and Dolly, considering the small size of the training sets, we train the models only for 100 steps with an effective batch size of 8. For OLMo2-SFT-Mix tasks, we train the model for 1,000 steps. We use HuggingFace Transformers library for training and VLLM library for efficient inference. Due to the computational resource limitations, the statistics of forgetting are collected in a single run.
Dataset and licenses. MMLU, BBH, and the Pile are released under MIT license. Truthful QA, Dolma, Redpajama, OLMo models, OLMo2 models, Pythia models, and MPT models are released under Apache 2.0 license. Tulu V2, OLMo2-Mix, and OLMo2-SFT-Mix are released under ODC-By license. Dolly is released under CC BY-SA 3.0 license.
Computational Infrastructure. We used 4 Quadro RTX A6000 GPUs for fine-tuning LLMs, and used 1 Quadrio RTX A6000 GPU for LLM inference.
@@
Figure 7: A side-by-side comparison between the matrices of forgetting, inner products of gradients and weight differences ($z_{ij}^{\textrm{g-w}}$), and the negative inner products of gradients ($z_{ij}^{\textrm{g-g}}$) we examined in Sec. 3.3.
@@
Figure 8: Visualized matrices of associations between learned tasks and forgotten examples on FLAN after fine-tuning OLMo-7B-Instruct, measured with binary exact match drop. Each colored pixel ($z_{ij}=1$) indicates forgetting of an upstream example $x_{j}$ after fine-tuning the model on the task $T_{i}$.
Appendix D Details of Forgetting Prediction and Replay
@@
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 12 and 13. The train-test split for the in-domain tasks is randomly generated.
@@
Training and evaluation details. We use Surprise Library 1.1.3222https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, MF, and KNN prediction models. For MF, we set the dimension of the learnable features (rank) as 5. We train the regression models for 1,000 epochs over the association matrices.
@@
For in-domain test splits, we randomly sample 30 upstream examples and assume ground truth forgetting is known for these examples. This is required for predicting forgetting on the rest of upstream examples by additive linear, MF, and KNN methods. We repeat the experiment 10 times.
@@
We used all-distilroberta-v1 sentence transformer model as the encoder in the implementation of the embedding similarity based prediction model of forgetting. At inference, given an upstream example $x_{j}$, we compute the averaged dot-product with all examples in the learned task $T_{i}$. We note that at inference time the approach does not require ground truth forgetting of a small number of examples. For a fair comparison with other matrix completion methods, we replace the prediction of the approach with ground truth forgetting on these examples.
Table 5: RMSE of forgetting prediction with matrix factorization (MF) models under different ranks.
Model / Rank 1 3 5 8 10
OLMo-7B 7.31 7.17 7.14 7.12 7.14
OLMo-7B-OOD 6.09 5.78 5.76 5.77 5.77
OLMo-1B 2.85 2.82 2.80 2.78 2.78
OLMo-1B-OOD 2.86 2.83 2.82 2.82 2.82
Sensitivity analysis to the rank in MF models. Table 5 summarizes the RMSE of forgetting prediction with MF models across different ranks $r$. The performance increases with $r$ in the beginning and then drops, indicating an overfit.
(a) OLMo-7B, full association matrix
(b) OLMo-7B, compontent $k=1$
(c) OLMo-7B, component $k=2$
(d) OLMo-7B, component $k=3$
(e) OLMo-7B, component $k=4$
@@
Figure 9: Reconstruction of $Z$ in OLMo-7B experiments with $k$-th singular value and vectors. Components of higher values of $k$ capture finer-grained details in $Z$.
(a) OLMo-7B-Instruct, full association matrix
(b) OLMo-7B-Instruct, compontent $k=1$
(c) OLMo-7B-Instruct, component $k=2$
(d) OLMo-7B-Instruct, component $k=3$
(e) OLMo-7B-Instruct, component $k=4$
Figure 10: Reconstruction of log-perplexity measured forgetting association matrix $Z$ in OLMo-7B-Instruct experiments with $k$-th singular value and vectors. Higher values of $k$ capture finer-grained details in $Z$.
@@
Replaying upstream examples in fine-tuning. We sparsely replay 1 mini-batch of 8 upstream examples every 32 steps of model updates while fine-tuning on new tasks. An exception is fine-tuning of OLMo-7B-Instruct models on Dolly, where we perform a replay every 8 steps given the smaller number of model training steps. Given predicted or ground truth forgetting $z_{i,1..N}$ on upstream examples $x_{1..N}$ when learning a new task $T_{i}$, we sample upstream examples to replay from a categorical distribution where $p(x_{j})\propto\exp(z_{i,j}/\tau)$, where $\tau$ is a temperature hyperparameter set as 0.1. The hyperparameter $\tau$ is tuned on a single validation task while reweighting replay examples with the ground truth forgetting $Z$.
Appendix E Details of the Example Similarity Metrics
@@
Textual similarity. We calculate the textual cosine similarity between learned tasks and upstream examples using TF-IDF vectorized representations of training examples in each learned task $T_{i}$ and an upstream example $x_{j}$. We also measure text representation similarity with final layer representations of OLMo-1B.
Inner products between projected gradients and model weight updates. The increase of the log perplexity ${z_{ij}}$ can be approximated with inner products $z^{\textrm{g-w}}_{ij}=\langle\nabla_{\theta}f(x_{j}),\theta_{T_{i}}-\theta_{0}\rangle$ under first-order Taylor expansion (Lee et al., 2019; Doan et al., 2020), where $\nabla_{\theta}f(x_{j})$ is the gradient of the loss of $x_{j}$ at the initial model before fine-tuning, and $\theta_{T_{i}}-\theta_{0}$ are the updates in the model weights after fine-tuning. Following Park et al. (2023); Xia et al. (2024), we use a random projection matrix $\bm{P}\sim\mathcal{N}_{|\theta|\times d}(0,1)$ to reduce the dimension of the gradients or the weight changes to save the cost of storing pre-computed statistics, which preserves the inner products with high probability (Johnson & Lindenstrauss, 1984).
@@
Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
Table 6: Validation log perplexity of the learned tasks after fine-tuning LMs with different replay strategies.
Model OLMo-1B OLMo-7B OLMo-7B OLMo-7B-Instruct
Dataset FLAN FLAN Tulu & Dolly OLMo2-SFT-Mix
No Replay 0.9723 0.7736 1.6397 0.6677
Random 0.9621 0.7691 1.6452 0.6671
MF-Offline 0.9601 0.7684 1.6294 0.6674
Targeted replay does not impede learning of new tasks, as shown with decreased validation perplexity. Table 6 summarizes log perplexity on the validation set of learned tasks after fine-tuning the models with different replay strategies. We notice that example replay and targeted replay often decrease validation perplexity of new tasks. This improvement of new task perplexity can be attributed to less forgetting of general knowledge during fine-tuning. Besides, the results imply that targeted replay does not simply trade off new task learning for reduced forgetting.
LLM leaderboard task evaluation. We evaluate downstream task performance of OLMo-7B-Instruct after fine-tuning on 4 held-out tasks from OLMo2-SFT-Mix (marked in Table 7). We evaluate on Open LLM Leaderboard tasks [^3] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about. For fine-tuned models, we compare no replay, replaying random examples, and replaying with forgetting predicted by offline MF. Tables 7 summarize the results.
We notice that fine-tuning OLMo-7B-Instruct over new tasks without any replay improves metrics scores on some tasks (MUSR, GPQA) but causes forgetting on other tasks (MMLU, BBH, IFEval). Among tasks where performance improved, we do not see benefits of example replay. Among tasks with decreased performance, we see offline-MF mitigates forgetting compared to no replay or random replay slightly.
We conjecture that replay-based approaches are not sufficient to significantly mitigate forgetting on their own, and can be combined with other approaches. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
Table 7: Average downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on 4 held-out OLMo2-SFT-Mix tasks.
@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
No Replay 17.21 36.76 21.04 40.20 27.99
Random 17.37 36.85 20.97 39.77 27.73
MF-Offline 17.43 36.89 21.14 39.50 27.69
@@
Learned Tasks Forgotten Domain Learned Tasks Forgotten Domain
@@
$k=2$ dolly/general_qa StackOverflow (Less) flan/mnli_mismatched None
@@
Appendix G Towards Interpreting Fine-Grained Associations
We visualize progressive reconstruction with $k$-th singular value and singular vectors for OLMo experiments in Figure 10.